# Near-Linear and Parameterized Approximations for Maximum Cliques in Disk Graphs

### v2

@@ Abstract
A disk graph is the intersection graph of (closed) disks in the plane. We consider the classic problem of finding a maximum clique in a disk graph. For general disk graphs, the complexity of this problem is still open, but for unit disk graphs, it is well known to be in P. The currently fastest algorithm runs in time $O(n^{7/3+o(1)})$, where $n$ denotes the number of disks [EspenantKM23, keil_et_al:LIPIcs.SoCG.2025.63]. Moreover, for the case of disk graphs with $t$ distinct radii, the problem has also recently been shown to be in XP. More specifically, it is solvable in time $O^{*}(n^{2t})$ [keil_et_al:LIPIcs.SoCG.2025.63]. In this paper, we present algorithms with improved running times by allowing for approximate solutions and by using randomization:
@@
    for disk graphs with $t$ distinct radii, we give a parameterized approximation scheme that, with a constant success probability, computes a $(1-\varepsilon)$-approximate maximum clique in expected time $\tilde{O}(f(t)\cdot(1/\varepsilon)^{O(t)}\cdot n)$.
@@ 1 Introduction
A disk graph is the intersection graph of closed disks in the plane, where the vertices are the disks, and two vertices are connected by an edge if and only if the two corresponding disks intersect. A unit disk graph is a disk graph in which all disks have the same radius, which can be assumed to be $1$. Disk graphs and unit disk graphs are probably the most extensively studied classes of geometric intersection graphs [Bonamy2021-ep, bonnet_et_al:LIPIcs.FSTTCS.2020.17, BreuK98, ClarkCJ90, Fishkin03, TkachenkoW25]. In particular, they are popular models of wireless communication networks, where each disk represents a wireless station, and the respective radii encode the transmission ranges. Theoretically, unit disk graphs and disk graphs are of special interest, because they generalize the familiar class of planar graphs. Indeed, the well-known circle-packing theorem (also called the the Koebe–Andreev–Thurston theorem [Koebe1936]) shows that every planar graph is a disk graph. Unlike planar graphs, disk graphs (or unit disk graphs) can be dense, with potentially $\Theta(n^{2})$ edges. Thus, algorithms on disk graphs often use an implicit representation (of the coordinates and radii of the disks) and derive edges on-demand.
Due to the special structure of disk graphs, it is an active research direction to study efficient approximation algorithms for classical NP-hard problems on disk graphs or unit disk graphs, such as maximum independent set [Chan2012-pk, Hunt1994-tu] and minimum dominating set [Gibson10algorithms, Nieberg2008-zi]. In this paper, we study the maximum clique problem on unit disk graphs and disk graphs. For a given graph $G=(V,E)$, the problem asks for a vertex set $S\subseteq V$ of maximum cardinality such that the induced subgraph on $S$ is a clique. The maximum clique problem on general graphs is a classical NP-hard problem and is also hard to approximate [Feige2002-va]. However, maximum clique can be solved in polynomial time on unit disk graphs and the problem has been actively studied recently. We first review prior work on this problem and then present our new results.
@@
For general disk graphs, the complexity of computing the maximum clique has been a long-standing open problem [AmbuhlW05, EspenantKM23, Fishkin03]. Recently, Keil and Mondal [keil_et_al:LIPIcs.SoCG.2025.63] considered the case where there are $t$ distinct disk radii and gave an algorithm that runs in polynomial time for every constant $t$. In particular, the running time is $O(2^{t}n^{2t}(f(n)+n^{2}))$, where $f(n)$ is the time to compute a maximum matching in an $n$-vertex bipartite graph. A maximum matching can be computed in time $\tilde{O}(m)$ using the near-linear time max-flow algorithm in the recent breakthrough in [Chen2022-sz], where $m$ is the number of edges in the bipartite graph, or in time $\tilde{O}(n^{2})$ by a combinatorial algorithm [Chuzhoy2024-lj]. A $2$-approximation can be easily found in linear time [Carmi2023-av] by computing a stabbing set of four points [AmbuhlW05]. Further, the problem admits an EPTAS [^2] EPTAS stands for efficient polynomial-time approximation scheme. with a running time of $2^{\tilde{O}(1/\varepsilon^{3})}n^{O(1)}$ and an exact sub-exponential time algorithm [Bonamy2021-ep] (both of which also generalize to unit ball graphs).
Finally, the problem has also been studied for intersection graphs of other classes of objects in the plane and has been shown to be NP-hard for rays [Cabello2013-yj], strings [Keil2022-op], ellipses and triangles [AmbuhlW05], and for sets of both rectangles and unit disks [bonnet_et_al:LIPIcs.FSTTCS.2020.17].
@@
For improving the currently best $O(n^{7/3+o(1)})$-time bound for unit disk graphs, any algorithm that follows the classical line of approach by Clark, Colbourn, and Johnson [ClarkCJ90] would effectively need to reduce the number of candidate disk centers to $o(n)$ or compute a maximum clique in the resulting co-bipartite graph in $O(n)$ time. Both of these options seem challenging. On the other hand, the algorithm by Tkachenko and Wang [TkachenkoW25] does not bring an improvement for instances with $\Omega(n)$-size cliques, and dense instances are in fact the hardest for the problem, see for example [Bonamy2021-ep].
@@
Our second result is for the general disk graph setting. As mentioned above, the EPTAS in [Bonamy2021-ep] runs in $2^{\tilde{O}(1/\varepsilon^{3})}\cdot O(n^{4})$ time while, for disk graphs with $t$ distinct radii, the algorithm by Keil and Mondal [keil_et_al:LIPIcs.SoCG.2025.63] runs in $O^{*}(n^{2t})$ time [^3] The standard notation $O^{*}(\cdot)$ hides polynomial factors.. This raises the question of whether such exponential dependencies w.r.t. $t$ (on $n$) and $\varepsilon$ can be improved or even removed altogether. We show that this is possible by considering parameterized $(1-\varepsilon)$-approximations w.r.t. $t$ and $\varepsilon$ and, as in the case of unit disk graphs, by using randomization.
@@
    For disk graphs with $t$ distinct radii, there is an efficient parameterized approximation scheme (EPAS) [^4] An efficient parameterized approximation scheme (EPAS), for some parameter $k$, is a $(1-\varepsilon)$-approximation algorithm that runs in $f(k,\varepsilon)n^{O(1)}$ expected time [FeldmannKEM20]. that computes with a constant success probability a $(1-\varepsilon)$-approximate maximum clique in $\tilde{O}(f(t)\cdot(1/\varepsilon)^{O(t)}\cdot n)$ time, for some (exponential) function $f(t)$; this result is presented in Section 4.
Our algorithm is based on the main idea behind the exact algorithm by Keil and Mondal. Suppose that for each of the radii one can guess the leftmost and rightmost disks in an optimal solution. Then, one can find two sets of disks $L$ and $R$ such that $L\cup R$ induces a co-bipartite graph and contains a maximum clique. The algorithm by Keil and Mondal [keil_et_al:LIPIcs.SoCG.2025.63] enumerates all possible choices to find the leftmost and rightmost disks for each radius. To speed up the running time for $(1-\varepsilon)$-approximation, we bring in three more ideas. First, we can assume w.l.o.g that for disks of radius $r_{i}$ that appear in the optimal clique, at least $\Theta(\varepsilon k^{*}/t)$ of them appear in the max clique where $k^{*}$ is the optimal clique size and can be approximated by a factor of $2$ in linear time – otherwise, we can skip disks of radius $r_{i}$ completely without losing more than $\varepsilon$ fraction. Second, for each radius $r_{i}$ we can randomly sample two disks $a_{i}$ and $b_{i}$ – if we repeat this enough times, one such pair is close to, in terms of rank, the leftmost and rightmost disks of radius $r_{i}$ in the optimal solution. Last, as in the case of unit disks, we compute a $(1-\varepsilon)$-approximate clique in the resulting co-bipartite graph using the algorithm from Section 2.
@@ 1.1 Definitions and notation
Let $a$ and $b$ be two points in the plane. The Euclidean distance between $a$ and $b$ is denoted by $\mathrm{dist}(a,b)$. The vertical slab of $a$ and $b$ is the region between the vertical lines through $a$ and $b$, that is, the set of all points whose $x$-coordinate lies between the $x$-coordinates of $a$ and $b$.
A graph is co-bipartite if its complement is bipartite. Let $X$ and $Y$ be sets of disks such that $G(X)$ and $G(Y)$ are both cliques. As observed in [ClarkCJ90], $G(X\cup Y)$ is co-bipartite. Let $A$ be a set. We denote its cardinality by $|A|$. Let $X^{*}\subseteq\mathcal{D}$ be a maximum-size clique in $G(\mathcal{D})$, i.e., a subset of disks that all pairwise intersect. A clique $X\subseteq\mathcal{D}$ of size at least $(1-\varepsilon)|X^{*}|$ is a $(1-\varepsilon)$-approximate (maximum) clique.
@@ 2 Approximating the Maximum Clique in a co-Bipartite Disk Graph
As explained by Agarwal, Efrat, and Sharir [AgarwalES00, Section 9.1], the problem reduces to maintaining the upper envelope of a set of $n$ totally defined continuous bivariate functions of constant description complexity in three dimensions, under insertions, deletions, and vertical ray shooting queries. The discussion of Agarwal, Efrat, and Sharir [AgarwalES00, Section 9.1] also shows that the complexity of these upper envelopes (and thus of the corresponding farthest-site Voronoi diagrams) is linear. Thus, the result follows from the data structure given in [DBLP:journals/dcg/KaplanMRSS20, Theorem 8.3] and its subsequent improvement [Liu22]. ∎
With the above data structure in hand, we can describe the algorithm for computing a $(1+\varepsilon)$-approximate matching in $\overline{G}$. As explained above, returning $n$ minus its cardinality gives us the cardinality of a $(1-\varepsilon)$-approximate clique in $G$.
@@
Let the given co-bipartite graph be $G(X\cup Y)$. We will call $X$ the left side, and $Y$ the right side. We recall that the algorithm of Hopcroft and Karp [HopcroftKarp73] proceeds in phases while maintaining the current matching $M$ (initially empty). In the beginning of each phase, we run a breadth-first search to partition the nodes of the graph into layers. We start the search from the unmatched nodes on the left side of the graph that form layer 0. On even layers of the search, we can traverse any unmatched edge. On odd layers of the search, we can only traverse a matched edge. The search terminates after the first layer containing an unmatched node on the right side of the graph (or when no further layers can be created, meaning that the current matching $M$ is a maximum matching). This gives us the length $k$ of a shortest augmenting path. Next, the algorithm finds a maximal set of node-disjoint augmenting paths of length $k$. This can be implemented by running a depth-first search from the unmatched nodes in the last layer. The depth-first search is only allowed to follow an edge leading to a yet unvisited node in the previous layer, and the paths in its tree must alternate between matched and unmatched edges, as in the breadth-first search.
@@
The above lemma allows us to compute the cardinality of the sought $(1-\varepsilon)$-approximate clique. Computing the corresponding subset of nodes is a bit more involved.
@@
Let $G$ be a co-bipartite disk graph with $n$ vertices. We can compute a $(1-\varepsilon)$-approximate clique of $G$ in expected $O((n/\varepsilon)\log^{4}n)$ time.

### v3

@@ Abstract
A disk graph is the intersection graph of (closed) disks in the plane. We consider the classic problem of finding a maximum clique in a disk graph. For general disk graphs, the complexity of this problem is still open, but for unit disk graphs, it is well known to be in P. The currently fastest algorithm runs in time $O(n^{7/3+o(1)})$, where $n$ denotes the number of disks [19, 29]. Moreover, for the case of disk graphs with $t$ distinct radii, the problem has also recently been shown to be in XP. More specifically, it is solvable in time $O^{*}(n^{2t})$ [29]. In this paper, we present algorithms with improved running times by allowing for approximate solutions and by using randomization:
@@
    for disk graphs with $t$ distinct radii, we give a parameterized approximation scheme that, with a constant success probability, computes a $(1-\varepsilon)$-approximate maximum clique in expected time $\tilde{O}(f(t)\cdot(1/\varepsilon)^{O(t)}\cdot n)$, for some (exponential) function $f(t)$.
@@ 1 Introduction
A disk graph is the intersection graph of closed disks in the plane, where the vertices are the disks, and two vertices are connected by an edge if and only if the two corresponding disks intersect. A unit disk graph is a disk graph in which all disks have the same radius, which can be assumed to be $1$. Disk graphs and unit disk graphs are probably the most extensively studied classes of geometric intersection graphs [5, 6, 7, 15, 22, 35]. In particular, they are popular models of wireless communication networks, where each disk represents a wireless station, and the respective radii encode the transmission ranges. Theoretically, unit disk graphs and disk graphs are of special interest, because they generalize the familiar class of planar graphs. Indeed, the well-known circle-packing theorem (also called the the Koebe–Andreev–Thurston theorem [30]) shows that every planar graph is a disk graph. Unlike planar graphs, disk graphs (or unit disk graphs) can be dense, with potentially $\Theta(n^{2})$ edges. Thus, algorithms on disk graphs often use an implicit representation (of the coordinates and radii of the disks) and derive edges on-demand.
Due to the special structure of disk graphs, it is an active research direction to study efficient approximation algorithms for classical NP-hard problems on disk graphs or unit disk graphs, such as maximum independent set [11, 25] and minimum dominating set [23, 32]. In this paper, we study the maximum clique problem on unit disk graphs and disk graphs. For a given graph $G=(V,E)$, the problem asks for a vertex set $S\subseteq V$ of maximum cardinality such that the induced subgraph on $S$ is a clique. The maximum clique problem on general graphs is a classical NP-hard problem and is also hard to approximate [20]. However, maximum clique can be solved in polynomial time on unit disk graphs and the problem has been actively studied recently. We first review prior work on this problem and then present our new results.
@@
For general disk graphs, the complexity of computing the maximum clique has been a long-standing open problem [3, 19, 22]. Recently, Keil and Mondal [29] considered the case where there are $t$ distinct disk radii and gave an algorithm that runs in polynomial time for every constant $t$. In particular, the running time is $O(2^{t}n^{2t}(f(n)+n^{2}))$, where $f(n)$ is the time to compute a maximum matching in an $n$-vertex bipartite graph. A maximum matching in a bipartite graph can be computed in time $\tilde{O}(m)$, where $m$ is the number of edges in the graph, using the near-linear time max-flow algorithm in the recent breakthrough in [13] with standard reductions to bipartite matching, e.g., the $O(m\log m)$-time fractional flow rounding technique in [26]; alternatively, it can be computed in time $\tilde{O}(n^{2})$ by a combinatorial algorithm [14]. Further, the problem admits a randomized EPTAS [^2] EPTAS stands for efficient polynomial-time approximation scheme. with a running time of $2^{\tilde{O}(1/\varepsilon^{3})}n^{O(1)}$ and an exact sub-exponential time algorithm [5].
Finally, the problem has also been studied for intersection graphs of other classes of objects in the plane and has been shown to be NP-hard for rays [9], strings [28], ellipses and triangles [3], and for sets of both rectangles and unit disks [6].
@@
For improving the currently best $O(n^{7/3+o(1)})$-time bound for unit disk graphs, any algorithm that follows the classical line of approach by Clark, Colbourn, and Johnson [15] would effectively need to reduce the number of candidate disk centers to $o(n)$ or compute a maximum matching in the resulting complement bipartite graph in $o(n^{4/3})$ time. Both of these options seem challenging. On the other hand, the algorithm by Tkachenko and Wang [35] does not bring an improvement for instances with $\Omega(n)$-size cliques, and dense instances are in fact the hardest for the problem, see for example [5].
@@
Our second result is for the general disk graph setting. As mentioned above, the EPTAS in [5] runs in $2^{\tilde{O}(1/\varepsilon^{3})}\cdot O(n^{4})$ time while, for disk graphs with $t$ distinct radii, the algorithm by Keil and Mondal [29] runs in $O^{*}(n^{2t})$ time [^3] The standard notation $O^{*}(\cdot)$ hides polynomial factors.. This raises the question of whether such exponential dependencies w.r.t. $t$ (on $n$) and $\varepsilon$ can be improved or even removed altogether. We show that this is possible by considering parameterized $(1-\varepsilon)$-approximations w.r.t. $t$ and $\varepsilon$ and, as in the case of unit disk graphs, by using randomization.
@@
    For disk graphs with $t$ distinct radii, there is an efficient parameterized approximation scheme (EPAS) [^4] An efficient parameterized approximation scheme (EPAS), for some parameter $k$, is a $(1-\varepsilon)$-approximation algorithm that runs in $f(k,\varepsilon)n^{O(1)}$ expected time [21]. that computes with a constant success probability a $(1-\varepsilon)$-approximate maximum clique in $\tilde{O}(f(t)\cdot(1/\varepsilon)^{O(t)}\cdot n)$ time, for some (exponential) function $f(t)$; this result is presented in Section 4.
Our algorithm is based on the main idea behind the exact algorithm by Keil and Mondal. Suppose that for each of the radii one can guess the leftmost and rightmost disks in an optimal solution. Then, one can find two sets of disks $L$ and $R$ such that $L\cup R$ induces a co-bipartite graph and contains a maximum clique. The algorithm by Keil and Mondal [29] enumerates all possible choices to find the leftmost and rightmost disks for each radius. To speed up the running time for $(1-\varepsilon)$-approximation, we bring in three more ideas. First, we can assume w.l.o.g that for disks of radius $r_{i}$ that appear in the optimal clique, at least $\Theta(\varepsilon k^{*}/t)$ of them appear in the max clique where $k^{*}$ is the optimal clique size, which can be approximated by a factor of $1/5$ in $O(n\log n)$ time (see section˜4.2) – otherwise, we can skip disks of radius $r_{i}$ completely without losing more than $\varepsilon$ fraction. Second, for each radius $r_{i}$ we can randomly sample two disks $a_{i}$ and $b_{i}$ – if we repeat this enough times, one such pair is close to, in terms of rank, the leftmost and rightmost disks of radius $r_{i}$ in the optimal solution. Last, as in the case of unit disks, we compute a $(1-\varepsilon)$-approximate clique in the resulting co-bipartite graph using the algorithm from Section 2.
@@ 1.1 Definitions and notation
Let $a$ and $b$ be two points in the plane. The Euclidean distance between $a$ and $b$ is denoted by $\mathrm{dist}(a,b)$. The vertical slab of $a$ and $b$ is the region between the vertical lines through $a$ and $b$, that is, the set of all points whose $x$-coordinates lie between the $x$-coordinates of $a$ and $b$.
A graph is co-bipartite if its complement is bipartite. Let $X$ and $Y$ be sets of disks such that $G(X)$ and $G(Y)$ are both cliques. As observed in [15], $G(X\cup Y)$ is co-bipartite. Let $A$ be a set. We denote its cardinality by $|A|$. Let $X^{*}\subseteq\mathcal{D}$ be a maximum-size clique in $G(\mathcal{D})$, i.e., a subset of disks that all pairwise intersect. A clique $X\subseteq\mathcal{D}$ of size at least $(1-\varepsilon)|X^{*}|$ is a $(1-\varepsilon)$-approximate (maximum) clique.
@@ 2 Approximating the Maximum Clique in a co-Bipartite Disk Graph
As explained by Agarwal, Efrat, and Sharir [2, Section 9.1], the problem reduces to maintaining the upper envelope of a set of $n$ totally defined continuous bivariate functions of constant description complexity in three dimensions, under insertions, deletions, and vertical ray shooting queries. The discussion of Agarwal, Efrat, and Sharir [2, Section 9.1] also shows that the complexity of these upper envelopes (and thus of the corresponding farthest-site Voronoi diagrams) is linear. Thus, the result follows from the data structure given in [27, Theorem 8.3] and its subsequent improvement [31]. ∎
With the above data structure in hand, we can describe the algorithm for computing a $(1-\varepsilon)$-approximate matching in $\overline{G}$. As explained above, returning $n$ minus its cardinality gives us the cardinality of a $(1-\varepsilon)$-approximate clique in $G$.
@@
Let the given co-bipartite graph be $G(X\cup Y)$. We will call $X$ the left side, and $Y$ the right side. We recall that the algorithm of Hopcroft and Karp [24] proceeds in phases while maintaining the current matching $M$ (initially empty). In the beginning of each phase, we run a breadth-first search to partition the nodes of the graph into layers. We start the search from the unmatched nodes on the left side of the graph that form layer 0. On even layers of the search, we can traverse any unmatched edge. On odd layers of the search, we can only traverse a matched edge. The search terminates after the first layer containing an unmatched node on the right side of the graph (or when no further layers can be created, meaning that the current matching $M$ is a maximum matching). This gives us the length $k$ of a shortest augmenting path. Next, the algorithm finds a maximal set of node-disjoint augmenting paths of length $k$. This can be implemented by running a depth-first search from the unmatched nodes in the last layer. The depth-first search is only allowed to follow an edge leading to a yet unvisited node in the previous layer, and the paths in its tree must alternate between matched and unmatched edges, as in the breadth-first search.
@@
For the case of unit disk graphs, Lemma˜3 can be improved. For this, we just need a data structure that plays the role of the one in Lemma˜2 and that is optimized for unit disk graphs. The following lemma shows how to do this, adapting a method by Efrat, Itai, and Katz [17] to our setting:
Lemma 4.
Let $S$ be a set of $n$ unit disks. Then, we can preprocess $S$ in $O(n\log n)$ time into a data structure that supports the following operations:
  • $\text{find}(q)$: given $q\in\mathbb{R}^{2}$, find a disk in the data structure that does not contain $q$, or report that no such disk exists (i.e., $q$ lies in the intersection of all the disks in the data structure); and
  • $\text{delete}(s)$: delete the unit disk $s$ from the data structure.
The time for a find-operation is $O(\log n)$, and the total time to delete all the disks from the data structure is $O(n\log n)$.
Proof sketch.
Our data structure is very similar to the structure by Efrat, Itai, and Katz [17, Section 5.1], and we refer the reader to their paper for further details.
Let $\mathcal{G}$ be a regular grid with cells of edge length $1/2$. During the preprocessing phase, we locate for every disk center in $S$ the cell of $\mathcal{G}$ that contains it. Let $C$ be a nonemtpy grid cell, and let $S_{C}$ be the unit disks whose centers lie in $C$. Since the diameter of $C$ is less than 11, the disks in $S_{C}$ have non-empty intersection (e.g., the center of $C$ must lie in all the disks of $S_{C}$). We compute the intersection $P_{C}$ of $S_{C}$. The intersection $P_{C}$ is a convex polygon-like structure whose boundary consists of arcs from $S_{C}$, where every disk from $S_{C}$ contributes at most one arc. The arcs appear along the upper and the lower boundary of $P_{C}$ in reverse order of the $x$-coordinates of the corresponding disk centers. Thus, the total complexity of $P_{C}$ is $O(|S_{C}|)$, and it can be computed in time $O(|S_{C}|\log|S_{C}|)$, using a divide-and-conquer strategy. As in the data structure by Efrat, Itai, and Katz, we compute for each $P_{C}$ two binary trees that represent the structure of the upper and the lower boundary of $P_{C}$, respectively, as in the well-known data structure by Overmars and van Leeuwen [33]. In these trees, the leaves correspond to the disks from $S_{C}$, in reverse $x$-order of their centers. The inner nodes represent the upper (or the lower) boundary of the intersection of the disks in the corresponding subtree (see the paper by Efrat, Itai, and Katz for the details). The binary trees can be computed in $O(n\log n)$ time, and the total time for the preprocessing phase is thus $O(n\log n)$.
A find-query can be answered as follows: given $q\in\mathbb{R}^{2}$, we first check if there is a non-empty grid cell $C$ that is far enough from $q$ such that $q$ must lie outside of $P_{C}$. If such a cell $C$ exists, we report an arbitrary disk from $P_{C}$ and are done (this takes $O(\log n)$ time, assuming that the nonempty grid cells are stored in a suitable data structure). Otherwise, there is a constant number of remaining candidate cells, and for each such candidate cell $C$, we determine whether $q$ lies outside of $P_{C}$. This can be done using the tree structures for $C$ to determine whether $q$ lies above or below the upper and the lower boundary of $P_{C}$. If $q$ lies outside of $P_{C}$, we can use the result of the tree-searches to obtain a disk in $S_{C}$ that does not contain $q$, as desired. Again, this takes $O(\log n)$ time.
To delete a a disk $s$ from $S$, we find the cell $C$ that contains the center of $s$, and we delete $s$ from $S_{C}$. For this, we must update the tree structures that represent the intersection $P_{C}$. The details for the algorithm and its analysis are given by Efrat, Itai, and Katz, resulting in the claimed running time. ∎
Using Lemma˜4, we get the following improved version of Lemma˜3 for unit disk graphs:
Lemma 5.
Let $G$ be a co-bipartite unit disk graph on $n$ vertices. A $(1-\epsilon)$-approximate matching of $\overline{G}$ can be computed in expected $O((n/\varepsilon)\log n)$ time.
Proof.
The proof is identical to the proof of Lemma˜3, but using the data structure from Lemma˜4 instead of Lemma˜2. ∎
Lemma˜3 and Lemma˜5 allow us to compute the cardinality of the sought $(1-\varepsilon)$-approximate clique. Computing the corresponding subset of nodes is a bit more involved.
@@
Let $G$ be a co-bipartite disk graph with $n$ vertices. We can compute a $(1-\varepsilon)$-approximate clique of $G$ in expected $O((n/\varepsilon)\log^{4}n)$ time. When $G$ is a unit disk graph, such a clique can be computed in expected $O((n/\varepsilon)\log n)$ time.
@@
We provide the analysis for general disk graphs. The case of unit disk graphs is handled in the same way, using the faster data structure from Lemma˜4.